# DRL-ORA: Distributional Reinforcement Learning with Online Risk Adaption

### v4

@@ Appendix A Appendix
where $N$ and $N^{\prime}$ denote the respective number of i.i.d. samples $q_{i},\,q_{j^{\prime}}\sim U([0,\,1])$ used to estimate the loss, and $h_{\tau}^{\kappa}$ is the quantile estimate generated by the Huber loss function (see (Huber 1992)) with a hyper-parameter threshold $\kappa$ to smooth the non-differentiable point at zero; that is,
@@
where $\delta_{t}$ is the TD-error, $\mathbb{I}\{\cdot\}$ is the indicator function, and normally $\kappa=1$ is chosen (see (Dabney et al. 2018a)). A corresponding sample-based risk-aware policy is then obtained by approximating $Q_{\beta}$ through the average of distorted return distribution samples. Based on the above settings, we are able to estimate the return distribution through a neural network structure, update the network parameters by by minimizing the IQN loss, and implement a risk-aware policy.
Proofs for Technical Results
Proof of Theorem 4: Based on Lemma 4.1 of (Cesa-Bianchi and Lugosi 2006) and (Agarwal, Gonen, and Hazan 2019), we can instead analyze a slightly different algorithm which draws only a single noise vector $\sigma\sim\exp(\eta)$, rather than drawing a fresh noise vector on every round. Proving regret bounds for this variant translates into asymptotically equivalent (expected) regret bounds for Algorithm 1. Set $l_{0}=-\sigma\cdot\alpha$. We have the following several lemmas and proposition hold which are used to prove Theorem 4.
@@
The proof is inspired by the proof of Lemma 2.1 in Shalev-Shwartz et al. (2012). Subtracting $\sum_{t}l_{t}(\alpha_{t}^{\prime})$ from both sides of the inequality and rearranging, the desired inequality can be rewritten as
@@
The proof is inspired by the proof of Theorem 1 in Agarwal, Gonen, and Hazan (2019). For any two functions $f_{1},\,f_{2}:\mathcal{A}^{\prime}\rightarrow\mathbb{R}$ and scalar $\sigma_{1},\,\sigma_{2}\in\mathbb{R}$, let
@@
$\min_{\boldsymbol{c},b\geq 0}$~~ $\frac{1}{K}\sum_{k=1}^{K}c_{k}$
@@
if $\sum_{k\in[K]/[K^{\prime}]}(Q^{\theta_{k}}_{t-1}-\tau_{t})\geq\sum_{k\in[K^{\prime}]}(\tau_{t}-Q^{\theta_{k}}_{t-1})$then
@@
The most computational consuming step in Algorithm is the sorting $b$-intercepts in the set $K^{\prime}]$ which has complexity $O(K\log(K))$, which much lower than that of the best known algorithm for linear program $O(K^{2.5})$ (see (Vaidya 1989)).
Supplementary Materials for Experiments
@@
Table 1 records, for ORA, the monitored LTV to the median of epistemic uncertainty distribution obtained in each period of the three experiments, as well as the corresponding risk parameters in aligned training periods. The hyper-parameter settings for CartPle, Atari games, Knapsack and Nano Drone Navigation tasks are listed in Table 3, 4 and 5. Parameters not listed are as default in IQN. The implementation and hyper-parameter settings of ART and TOP in Atari games and Knapsack environments follows the open source (Moskovitz et al. 2021; Liu, van Kampen, and De Croon 2023). The experiment was conducted on a Windows platform with an Intel i5-13400F and an Nvidia RTX 5060. The dependencies and libraries required for each task can be found in the code appendix.
@@
Table 3: Hyper-parameters in CartPle and Atari game environments

### v5

@@ Appendix A Preliminaries on the IQN Loss Function
where $N$ and $N^{\prime}$ denote the respective number of i.i.d. samples $q_{i},\,q_{j^{\prime}}\sim U([0,\,1])$ used to estimate the loss, and $h_{\tau}^{\kappa}$ is the quantile estimate generated by the Huber loss function (see huber1992robust) with a hyper-parameter threshold $\kappa$ to smooth the non-differentiable point at zero; that is,
@@
where $\delta_{t}$ is the TD-error, $\mathbb{I}\{\cdot\}$ is the indicator function, and normally $\kappa=1$ is chosen (see dabney2018implicit). A corresponding sample-based risk-aware policy is then obtained by approximating $Q_{\beta}$ through the average of distorted return distribution samples. Based on the above settings, we are able to estimate the return distribution through a neural network structure, update the network parameters by by minimizing the IQN loss, and implement a risk-aware policy.
Appendix B Proofs for Technical Results
Proof of Theorem 4: Based on Lemma 4.1 of cesa2006prediction and agarwal2019learning, we can instead analyze a slightly different algorithm which draws only a single noise vector $\sigma\sim\exp(\eta)$, rather than drawing a fresh noise vector on every round. Proving regret bounds for this variant translates into asymptotically equivalent (expected) regret bounds for Algorithm 1. Set $l_{0}=-\sigma\cdot\alpha$. We have the following several lemmas and proposition hold which are used to prove Theorem 4.
@@ Appendix B Proofs for Technical Results
The proof is inspired by the proof of Lemma 2.1 in shalev2012online. Subtracting $\sum_{t}l_{t}(\alpha_{t}^{\prime})$ from both sides of the inequality and rearranging, the desired inequality can be rewritten as
@@
The proof is inspired by the proof of Theorem 1 in agarwal2019learning. For any two functions $f_{1},\,f_{2}:\mathcal{A}^{\prime}\rightarrow\mathbb{R}$ and scalar $\sigma_{1},\,\sigma_{2}\in\mathbb{R}$, let
@@
$\min_{\boldsymbol{c},b\geq 0}\penalty 10000\ \penalty 10000\$$\frac{1}{K}\sum_{k=1}^{K}c_{k}$
@@
if $\sum_{k\in[K]/[K^{\prime}]}(Q^{\theta_{k}}_{t-1}-\tau_{t})\geq\sum_{k\in[K^{\prime}]}(\tau_{t}-Q^{\theta_{k}}_{t-1})$, then
@@
The most computational consuming step in Algorithm is the sorting $b$-intercepts in the set $K^{\prime}]$ which has complexity $O(K\log(K))$, which much lower than that of the best known algorithm for linear program $O(K^{2.5})$ (see vaidya1989speeding).
Appendix C Supplementary Materials for Experiments
Figure 6: Impact of ensemble size on ORA’s training performance and computational cost in CartPole environment. Results averaged over 3 runs per configuration (N = 2, 4, 8, 16, 32). Error bars show standard deviation. The time and memory overheads of IQN and TOP are indicated on the left and right sides of the second subplot, respectively.
Figure 7: ORA with 90% confidence interval.
@@ Appendix C Supplementary Materials for Experiments
Computational Test
Although the ensemble network is inseparable from our method, making it impossible to investigate the ablation improvement of the ensemble network, the test on ensemble size $K$ demonstrates that the ensemble network itself does not exhibit a significant correlation with performance improvement (Figure 6). Further computational tests reveal that the time overhead grows only slightly with the ensemble size $K$, while the associated memory overhead is negligible. This demonstrates that introducing the ensemble network does not incur a substantial computational burden. As a reference, we have also included the average time and space consumption of the TOP and IQN methods (the ensemble size is not applicable). The comparison further confirms that our method only requires slightly higher memory usage, and with the default $K$, the additional computational time is insignificant. Note that aforementioned experiments are conducted on the same GPU-accelerated device. Despite variations in model architectures of these approaches and consequently their hyperparameter configurations (which are specified in the following subsection), a uniform batch size of 64 was employed for all methods, with each using the parameter set determined to be optimal for it.
Table 2 records ORA’s left-truncated variance (LTV), risk parameters, and median epistemic uncertainty per training period for all three experiments. The experimental results of ORA (with distortion function CVaR and quantile) and baselines are shown in Figure 8. Across three diverse Atari environments, our ORA method demonstrates statistically significant dominance ($p$ < 0.001), achieving massive to large Rank-Biserial effect sizes against the TOP baseline (0.996 in Hero, 0.875 in MsPacman, 0.566 in SpaceInvaders) and substantial effects against ART (0.672, 0.511, 0.388, respectively). For these three environments, the discretization of $\mathcal{A}^{\prime}$ follows a grid size accuracy of 5, while in Nano Drone and knapsack, the grid size is bucketed to 1. The perturbation parameter $\eta$ for the FTPL algorithm is set to 0.5 in all experiments. The hyper-parameter settings for CartPole, Atari games, Knapsack and Nano Drone Navigation tasks are listed in Table 3, 4 and 5. Parameters not listed are as default in IQN. The implementation and hyper-parameter settings of ART and TOP in Atari games and Knapsack environments follows the open source [moskovitz2021tactical, liu2022adaptive]. The experiment was conducted on a Windows platform with an Intel i5-13400F and an Nvidia RTX 5060. The dependencies and libraries required for each task can be found in the code appendix.
@@
Table 3: Hyper-parameters in CartPole and Atari game environments